P500,000 was deposited at an interest of 6% compounded quarterly. Compute the compound interest after 4 years and 9 months.


Answer: P163,475.37


1
Expert's answer
2022-03-21T18:18:02-0400

4 years 9 months = 19 quarters.

The interest per quarter = "\frac{0.06}{4}=0.015"


The total amount after 4 years 9 months:

"y=a(1+r)^t=500000\times(1+0.015)^{19}=663475.37"


The total interest = 663475.37 - 500000 = P163475.37


Answer: P163475.37